OpenStudy (anonymous):

Factor completely: -12x4y - 9x3y2 + 3x2y3 Answer 3x2y(-4x2 + 3xy - y2) -3x2y(4x2 + 3xy - y2) -3x2y(-4x2 - 3xy + y2) -3(4x2y + 3x3y2 - x2y3)

OpenStudy (anonymous):

\[6 y^3-27 y^2-48 y=3 y \left(2 y^2-9 y-16\right) \]

OpenStudy (anonymous):

\[3\left( x ^{2}y \right) \left( x + y \right) \left( y - 4x \right)\]